OpenStudy (anonymous):

Hi, I'm to find the integral of 40e^(0.34657t) dt Can someone please explain how I would do this?

OpenStudy (anonymous):

\[\int k e^{at}dt\] where k and a are constants \[k\int e^{at}dt\] \[k \frac{e^at}{a} + C\] C is integration constant

OpenStudy (anonymous):